OpenStudy (anonymous):

What is the vertex of the graph of y = –|2x + 4|+ 3? (–2, –3) (–2, 3) (2, –3) (2, 3)

OpenStudy (anonymous):

can you help me pretty please ?

OpenStudy (phi):

you want the x value that makes the "stuff" inside the absolute value | | be zero when that happens, the y value will be 3

OpenStudy (anonymous):

i dont understand ?

OpenStudy (anonymous):

can you explain

OpenStudy (phi):

First, do you know what they mean by *vertex* ? if you plot this equation y = –|2x + 4|+ 3 the vertex is where it "changes direction" see http://www.wolframalpha.com/input/?i=plot+y+%3D+ –%7C2x+%2B+4%7C%2B+3

OpenStudy (phi):

It changes direction at the spot where –|2x + 4| is zero in other words, where 2x+4 is 0 you could write this as 2x+4 =0 and "solve" for x add -4 to both sides 2x +4 -4 = -4 the +4 and -4 add up to 0, so this is the same as 2x + 0 = -4 adding 0 to something does not change it, so this is 2x = -4 divide both sides by 2 x = -4/2 that simplifies to x= -2

OpenStudy (anonymous):

Now i understand which one it be (–2, –3) (–2, 3) (2, –3) (2, 3)

OpenStudy (anonymous):

??

OpenStudy (phi):

Now i understand good

OpenStudy (anonymous):

what would the anser be

OpenStudy (anonymous):

answer